OpenStudy (anonymous):

Which of the following is a simplification of the expression [(3*4)+(1*5)]*37/51

OpenStudy (anonymous):

ok, wat's the first thing that u did? :)

OpenStudy (anonymous):

3*4

OpenStudy (anonymous):

gud :) wat next? and this time show me the entire line :)

OpenStudy (anonymous):

(3*4+(1*5)= 17 * 37/51=

OpenStudy (anonymous):

629

OpenStudy (anonymous):

u still need to divide 629 by 51 :)

OpenStudy (goformit100):



OpenStudy (anonymous):

12?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

3x4= 12 yes but she then added 1x5=5 therefore 12+5=17

OpenStudy (anonymous):

@goformit100

OpenStudy (anonymous):

629/51 = 12.33 , that is ur answer :)